# Causal Graph Neural Networks for Healthcare

### v2

@@ 3 Methodological advances
The translation of causal theory into practical healthcare applications requires architectures that address specific clinical capabilities. We organise methodological advances not by technical taxonomy but by the clinical problems they solve, such as disentangling genuine causal mechanisms from confounding, predicting interventional effects without experimental data, generating patient-specific counterfactuals, achieving robust generalisation across heterogeneous settings, and ensuring fairness through causal rather than purely statistical criteria.
@@ 3.1 Disentangling causal signal from confounding
Biomedical datasets are rife with spurious correlations that mislead standard machine learning models. In brain imaging, global arousal states create correlations between all regions, in molecular networks, batch effects generate artificial associations between unrelated proteins, in clinical records, treatment selection bias produces correlations reflecting physician decision-making rather than therapeutic efficacy [26, 27]. The challenge is to decompose observed graphs into causal and spurious components, a problem addressed through information-theoretic disentanglement that maximises mutual information between causal features and outcomes and minimises spurious correlations [49].
@@
The DisC (Debiasing via Disentangled Causal Substructure) framework operationalises this separation through a dual-encoder architecture [42]. Rather than treating graphs monolithically, DisC maintains parallel GNN encoders: one trained to extract causal features that consistently predict outcomes across different environments, another capturing environment-specific spurious patterns. The contrastive learning objective enforces the independence between these representations, ensuring that causal and spurious features remain disentangled. For psychiatric diagnosis, this architectural separation was proven to be immensely effective. Traditional GNNs identified widespread hyperconnectivity in the cerebellar regions of depressed patients, a finding inconsistent with the clinical understanding that depression primarily influences affective rather than motor circuits. DisC’s causal decomposition correctly identified sparse cerebellar involvement while highlighting strong connections around the left rectus, a region known to be critical for treatment response to antidepressants [43]. This alignment with established neurobiology demonstrates that causal disentanglement provides mechanistically meaningful insights, not merely improved prediction.
@@
The CI-GNN (Causality-Inspired Graph Neural Network) advances this concept through Granger causality-inspired conditional mutual information, providing principled measures of causal strength for graph edges [44, 43]. The model identifies influential subgraphs that demonstrate genuine causal connections through disentangled subgraph-level representation learning. In autism spectrum disorder evaluations on the ABIDE dataset, CI-GNN revealed tight connections within the somatomotor network consistent with motor impairments characteristic of ASD, findings replicated across independent cohorts and suggesting genuine causal relationships rather than dataset-specific artefacts [44]. The clinical imperative was that clinicians needed to know which connections to target through neurostimulation, pharmacotherapy, or behavioural interventions, not merely that “connectivity is abnormal.” The IGCL-GNN framework extends these principles through multi-objective optimisation that simultaneously enhances causal features and suppresses environment-specific spurious patterns, recognising that robust clinical deployment requires balancing predictive performance on current datasets with generalisation to unseen environments [49].
@@ 3.2 Predicting interventional effects without experimental data
The CaT-GNN (Causal Temporal Graph Neural Network) implements interventional reasoning through architectural modifications encoding backdoor adjustment [51]. The framework first identifies which nodes represent causal factors versus environmental confounders, and then applies mixup augmentation specifically to environmental nodes. This selective augmentation breaks spurious correlations and preserves causal relationships, achieving robust prediction across different patient populations and clinical settings. For blood pressure monitoring, the CiGNN framework demonstrates the clinical utility of interventional prediction by first using the neural-accelerated causal discovery (FCI algorithm) to identify physiological signals causally related to blood pressure variation, then constructing spatiotemporal GNN architectures guided by the discovered causal structure [52, 53]. This causal feature selection reduces the dimensionality from 200+ potential features to 23 causal indicators while improving both accuracy and interpretability, identifying, for example, that specific heart rate variability patterns causally influence blood pressure through autonomic regulation and certain PPG morphological features affect pressure through arterial stiffness mechanisms.
The iVGAE (interventional Variational Graph Auto-Encoder) provides theoretical grounding for interventional GNNs by incorporating interventional layers in both encoder and decoder networks, ensuring that learnt representations explicitly encode causal rather than merely associational relationships [41]. When the model identifies a connection between proteins or brain regions, it learns whether this connection represents a causal pathway that would persist under intervention rather than a spurious correlation arising from confounding. For drug discovery, PDGrapher leverages protein-protein interaction networks as approximations of causal graphs, operating under causal sufficiency assumptions to predict combinatorial therapeutic targets [54]. This causally-inspired approach outperformed traditional methods in identifying effective perturbagens across nine cell lines with chemical perturbations, demonstrating that even imperfect causal structures, PPI networks contain both direct causal effects and indirect associations, provide superior guidance compared to purely correlational feature selection.
@@ 3.4 Achieving robustness across heterogeneous clinical settings
Healthcare AI deployment requires models that generalise across institutions with different patient demographics, clinical protocols, and measurement technologies, precisely the distribution shift conditions where correlation-based models fail catastrophically. The fundamental principle underlying invariant learning is that causal relationships remain stable across contexts while spurious correlations change [59]. In medical terms, the biological mechanisms by which drugs treat diseases should be consistent across populations, even though confounding factors such as demographics, comorbidities, and healthcare practices vary widely. This principle provides a powerful criterion for distinguishing causal from spurious patterns that hold across diverse environments and are likely causal.
The IGCL-GNN framework operationalises invariant causal learning through multi-objective optimisation that enhances causal features and suppresses environment-specific spurious patterns [49]. The key insight is that robust clinical deployment requires balancing competing objectives such as maximising predictive performance on current datasets and also ensuring generalisation to unseen environments. In multi-site brain imaging studies, CI-GNN maintained AUC scores of 0.81–0.84 across institutions while traditional GNNs varied dramatically from 0.68–0.85, demonstrating the value of learning causal invariances rather than site-specific correlations [43, 44]. This stability under distribution shift is not just a technical achievement, but a prerequisite for clinical trust—models that collapse when moved between hospitals cannot be deployed safely.
@@ 3.5 Ensuring causal fairness in clinical algorithms
The RaVSNet framework demonstrates the clinical impact of causal fairness correction through explicit modelling of prescription bias [63]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how historical prescription patterns, physician preferences, and institutional protocols influence medication choices independently of therapeutic benefit. The system generates causal effect matrices, adjusting the probabilities of recommendations to reflect genuine therapeutic value rather than the tradition of prescribing. In elderly patient populations, RaVSNet identified systematic under-prescription of effective medications to patients over 75 years of age, not due to legitimate contraindications, but to conservative prescribing practices. Correcting for these causal biases resulted in 18% improvement in treatment outcomes, demonstrating tangible equity gains [63]. This aligns with the disparate impact doctrine in anti-discrimination law, which prohibits practices creating group disparities unless justified by legitimate necessity. Causal decomposition provides quantitative evidence for whether less discriminatory alternatives exist, achieving equivalent clinical utility [35].
@@ 4.1.1 Mechanistic brain network analysis
Brain network analysis exemplifies the critical need for causal over correlational approaches. The human brain comprises approximately 86 billion neurons that form trillions of connections, creating one of the most complex networks in nature. Traditional graph-theoretical methods, while foundational, often fall short in capturing the high-dimensional and dynamic nature of brain connectivity [64]. The brain naturally forms a graph where regions of interest serve as nodes and functional connectivity defines edges, yet correlation-based analysis frequently identifies spurious connections driven by global brain states, arousal levels, or measurement artefacts rather than genuine neural communication pathways [43, 65].
@@
The BPI-GNN framework advances psychiatric diagnosis through prototype learning that identifies biologically meaningful subtypes of psychiatric disorders [66]. Beyond discriminating patients from healthy controls, the model revealed subtypes with distinct clinical profiles and treatment responses. In major depressive disorder, different subtypes exhibited distinct patterns of response to selective serotonin reuptake inhibitors versus serotonin-norepinephrine reuptake inhibitors, suggesting that network-based biomarkers could guide personalised treatment selection. These subtypes also corresponded to differences in gene expression profiles, providing biological validation of computationally identified clusters and suggesting potential molecular targets for subtype-specific interventions [66]. The model achieved superior performance against 11 popular brain network classification methods in three psychiatric datasets, demonstrating both technical excellence and clinical relevance.
For epileptic seizure prediction, the Dynamic Instance-level Graph Learning Network (DIGLN) captures patient-specific causal relationships between intracranial electroencephalogram channels that evolve over time [57]. By learning both the organisation of epileptic networks and their temporal evolution, DIGLN demonstrated the ability to evaluate multiple overlapping patterns of brain functional change, particularly in connectivity between the visual and sensorimotor domains [67]. This patient-specific approach achieved 91.7% sensitivity with only 0.18 false positives per hour, enabling reliable seizure warning systems that could transform the lives of epilepsy patients through timely intervention.
@@ 4.1.2 Multi-omics cancer subtyping
Causal methodology transforms cancer classification by focusing on causally related genes rather than on all differentially expressed features. The CautionGCN framework employs a causal multi-head autoencoder for causal feature selection, isolating characteristic genes with genuine causal relationships to cancer subtypes [69]. This approach addresses a fundamental problem in cancer genomics, where thousands of genes show differential expression, but only a subset causally drives tumorigenesis. By constructing causal gene-gene graphs that capture true driver relationships, MoCaGCN achieves a superior classification using only 89 genes compared to more than 1,000 in traditional approaches [70]. The Multi-omics Causal Graph Convolutional Network for lung adenocarcinoma classification incorporates causal structure learning algorithms combined with prior knowledge from gene-gene interaction databases, improving performance and generalisation by considering established biological interactions.
The identified causal genes provide more than classification accuracy. In lung adenocarcinoma, MoCaGCN identified KRAS-driven metabolic reprogramming as a causal mechanism of immune evasion, suggesting combination therapies targeting both KRAS and metabolic pathways [70]. These mechanistic insights have led to clinical trials that combine MEK inhibitors with immunotherapy, demonstrating how causal graph methods can accelerate translational research. The MOGONET framework takes advantage of both omics features and correlations among samples described by similarity networks, using the Correlation Discovery Network to explore cross-omics correlations in the label space [71]. This approach recognises that different omics layers may capture distinct aspects of cancer biology, with genomics revealing mutational drivers, transcriptomics showing pathway activation, and proteomics indicating functional consequences. By learning these multi-level representations jointly, MOGONET achieved superior performance in breast invasive carcinoma subtype classification while identifying layer-specific biomarkers.
@@ 4.2 Prognosis: predicting what interventions prevent
In a cohort of 3,847 patients with lung cancer surgery, KAMLN identified high-risk patients 72 hours earlier than standard risk scores by recognising early causal patterns [72]. This early warning enabled prophylactic interventions that reduced complication rates by 23% and shortened hospital stays by an average of 2.7 days. The interpretability of the model, showing not only risk, but also causal pathways, helped clinicians understand and trust the predictions, which proved crucial for clinical adoption. The explicit modelling of causal dependencies between complications has important equity implications beyond predictive performance. Patients in underserved communities, who often have more advanced disease states, benefit disproportionately from early identification of causal chains. In the cohort of lung cancer surgery, patients identified as high-risk 72 hours earlier by causal modelling received prophylactic interventions that reduced complication rates by 23%, with greater reductions observed in demographic subgroups historically experiencing higher complication rates [72]. This shows how causal reasoning can improve both accuracy and equity by enabling more proactive, mechanism-targeted clinical care.

### v3

@@ 3 Methodological advances
The translation of causal theory into practical healthcare applications requires architectures that address specific clinical capabilities. We organise methodological advances by the clinical capabilities they enable, disentangling genuine causal mechanisms from confounding, predicting interventional effects without experimental data, generating patient-specific counterfactuals, achieving robust generalisation across heterogeneous settings, and ensuring fairness through causal rather than purely statistical criteria, discussing within each capability the technical approaches that address it. From a complementary technical perspective, methods reviewed here rest on distinct theoretical foundations that determine their causal assumptions and evidentiary requirements. Intervention-invariance methods (Sections 3.1, 3.4) assume that causal features produce stable predictions across environments while spurious correlations shift with context; these require multi-environment data but not necessarily explicit causal graphs. Structural causal methods (Sections 3.2, 3.3) implement do-calculus through neural architectures, requiring explicit graph specification, causal sufficiency, and faithfulness assumptions; these provide formal identification guarantees but demand stronger prior knowledge. Knowledge-guided methods incorporate biomedical priors, protein-protein interactions, gene regulatory networks, anatomical atlases, as structural constraints without formal causal identification; these leverage domain expertise but inherit the epistemic status of their knowledge sources, which may conflate correlation with causation (Section 2.1). Table 2 maps individual methods to their specific assumptions within this taxonomy.
@@ 3.1 Disentangling causal signal from confounding
Biomedical datasets are rife with spurious correlations that mislead standard machine learning models. Standard GNNs, although naturally suited to graph-structured data, learn to exploit whatever statistical patterns maximise predictive performance regardless of causal validity or biological plausibility. In brain imaging, global arousal states create spurious correlations between all regions that GNNs readily exploit [57]; in molecular networks, batch effects create artificial associations between unrelated proteins [58]; in clinical data, treatment selection bias generates correlations between drug prescriptions and outcomes that reflect physician decision-making rather than therapeutic efficacy [26, 27]. The challenge is to decompose observed graphs into causal and spurious components, a problem addressed through information-theoretic disentanglement that maximises mutual information between causal features and outcomes and minimises spurious correlations [59].
@@
The DisC (Debiasing via Disentangled Causal Substructure) framework operationalises this separation through a dual-encoder architecture [52]. Rather than treating graphs monolithically, DisC maintains parallel GNN encoders: one trained to extract causal features that consistently predict outcomes across different environments, another capturing environment-specific spurious patterns. The contrastive learning objective enforces the independence between these representations, ensuring that causal and spurious features remain disentangled. For psychiatric diagnosis, this architectural separation was proven to be immensely effective. Traditional GNNs identified widespread hyperconnectivity in the cerebellar regions of depressed patients, a finding inconsistent with the clinical understanding that depression primarily influences affective rather than motor circuits. CI-GNN’s causal decomposition correctly identified sparse cerebellar involvement while highlighting strong connections around the left rectus, a region previously identified as critical for antidepressant treatment response prediction [53]. Evaluated on the REST-meta-MDD dataset (n=1,604), this alignment with established neurobiology demonstrates that causal disentanglement provides mechanistically meaningful insights, not merely improved prediction.
@@
The CI-GNN (Causality-Inspired Graph Neural Network) advances this concept through Granger causality-inspired conditional mutual information, providing principled measures of causal strength for graph edges [53, 52]. The model identifies influential subgraphs that demonstrate genuine causal connections through disentangled subgraph-level representation learning. In autism spectrum disorder evaluations on the ABIDE dataset, CI-GNN revealed tight connections within the somatomotor network consistent with motor impairments characteristic of ASD, findings replicated across independent cohorts and suggesting genuine causal relationships rather than dataset-specific artefacts [53]. The clinical imperative was that clinicians needed to know which connections to target through neurostimulation, pharmacotherapy, or behavioural interventions, not merely that ”connectivity is abnormal.” The IGCL-GNN framework extends these principles through multi-objective optimisation that simultaneously enhances causal features and suppresses environment-specific spurious patterns, recognising that robust clinical deployment requires balancing predictive performance on current datasets with generalisation to unseen environments [59].
Notably, these findings depend on the chosen parcellation resolution. CI-GNN operates on 116-node brain atlases where each node aggregates thousands of neurons; causal relationships at this mesoscale resolution reflect functional circuit organisation rather than individual neuronal connectivity, with direct implications for which neurostimulation or pharmacological interventions could target the identified causal pathways. Figure 3 illustrates these architectural principles applied to functional neuroimaging data. While the DisC and iVGAE frameworks were originally developed on synthetic benchmark datasets, their conceptual translation here to brain connectivity analysis could demonstrate how causal disentanglement and interventional prediction could address the specific confounding challenges in clinical neuroimaging, separating genuine neural circuit dysfunction from global arousal states, scanner artefacts, or site-specific acquisition protocols.
Figure 3: Canonical causal graph neural network architectures illustrated for conceptual brain connectivity analysis. While the underlying methods were developed on synthetic benchmark datasets, we illustrate their application to functional neuroimaging to demonstrate translation potential for clinical contexts. A, The DisC (Debiasing via Disentangled Causal Substructure) framework [52] achieves causal disentanglement through a dual-encoder architecture. An edge mask generator partitions input graphs into causal and spurious subgraphs, with separate GNN modules encoding each into disentangled representations. The loss comprises two components: $\mathcal{L}_{CE}$, a weighted cross-entropy loss that up-weights samples where the bias GNN struggles, directing the causal encoder toward genuinely predictive features; and $\mathcal{L}_{GCE}$, a generalised cross-entropy loss that amplifies gradients on easily-classified samples, directing the bias encoder toward environment-specific shortcuts. Causal features ($Z_{C.1}$$Z_{C.4}$) produce stable predictions across environments while spurious features ($Z_{S.1}$$Z_{S.3}$) are explicitly identified and discarded. B, The iVGAE (interventional Variational Graph Auto-Encoder) framework [45] implements Pearl’s do-operator through differentiable neural layers. The interventional encoder maps observed graphs to a causal latent space $Z$, while the interventional decoder incorporates do-layers that simulate graph surgery. The output $P(\mathbf{D}|\textit{do}(\mathbf{X}))$ represents the interventional distribution of downstream variables $\mathbf{D}$ given intervention $\textit{do}(\mathbf{X}=x)$, enabling prediction of effects that persist under manipulation rather than spurious correlations reflecting confounding. Applied to brain networks, such architectures could distinguish genuine neural communication pathways from global arousal-induced correlations, identifying intervention targets for neuromodulation or pharmacotherapy.
@@ 3.2 Predicting interventional effects without experimental data
The CaT-GNN (Causal Temporal Graph Neural Network) implements interventional reasoning through architectural modifications encoding backdoor adjustment [61]. The framework first identifies which nodes represent causal factors versus environmental confounders, and then applies mixup augmentation specifically to environmental nodes. This selective augmentation breaks spurious correlations and preserves causal relationships, achieving robust prediction across different patient populations and clinical settings. For blood pressure monitoring, the CiGNN framework demonstrates the clinical utility of interventional prediction by first using the neural-accelerated causal discovery (FCI algorithm) to identify physiological signals causally related to blood pressure variation, then constructing spatiotemporal GNN architectures guided by the discovered causal structure [51, 62]. This causal feature selection reduces dimensionality from 222 initial wearable features to a sparse set of direct and indirect causal indicators while improving both accuracy and interpretability, identifying, for example, that specific PPG morphological features (e.g., pulse arrival time, waveform characteristics) causally influence blood pressure through arterial stiffness mechanisms.
The iVGAE (interventional Variational Graph Auto-Encoder) provides theoretical grounding for interventional GNNs by incorporating interventional layers in both encoder and decoder networks, ensuring that learnt representations explicitly encode causal rather than merely associational relationships [45]. When the model identifies a connection between proteins or brain regions, it learns whether this connection represents a causal pathway that would persist under intervention rather than a spurious correlation arising from confounding. For drug discovery, PDGrapher leverages protein-protein interaction networks as approximations of causal graphs, operating under causal sufficiency assumptions to predict combinatorial therapeutic targets [50]. Evaluated on LINCS L1000 chemical perturbation data across nine cell lines (A549, MCF7, PC3, HT29, A375, and others), PDGrapher outperformed competing methods by 4.5–13.4% in detecting ground-truth targets, demonstrating that even imperfect causal structures, PPI networks contain both direct causal effects and indirect associations, provide superior guidance compared to purely correlational feature selection.
@@ 3.4 Achieving robustness across heterogeneous clinical settings
While disentanglement methods (Section 3.1) achieve robustness by explicitly separating causal from spurious representations, invariant learning approaches pursue the same objective through a complementary mechanism: optimisation constraints that enforce feature stability across environments without requiring explicit architectural separation. Disentanglement provides interpretable causal-spurious decomposition; invariant learning provides deployment guarantees with potentially simpler architectures. Some methods, such as IGCL-GNN, integrate both principles.
Healthcare AI deployment requires models that generalise across institutions with different patient demographics, clinical protocols, and measurement technologies, precisely the distribution shift conditions where correlation-based models fail catastrophically. The fundamental principle underlying invariant learning is that causal relationships remain stable across contexts while spurious correlations change (Figure 2B) [67]. In medical terms, the biological mechanisms by which drugs treat diseases should be consistent across populations, even though confounding factors such as demographics, comorbidities, and healthcare practices vary widely. This principle provides a powerful criterion for distinguishing causal from spurious patterns that hold across diverse environments and are likely causal.
The IGCL-GNN framework operationalises invariant causal learning through multi-objective optimisation that enhances causal features and suppresses environment-specific spurious patterns [59]. The key insight is that robust clinical deployment requires balancing competing objectives such as maximising predictive performance on current datasets and also ensuring generalisation to unseen environments. In multi-site brain imaging studies, CI-GNN achieved mean accuracy of 0.71–0.73 across 17 acquisition sites compared to 0.64–0.69 for baseline methods, with consistent improvements in 13 of 17 sites despite individual site accuracy ranging from 0.60 to 0.85, demonstrating the value of learning causal invariances rather than site-specific correlations [52, 53]. This stability under distribution shift is not just a technical achievement, but a prerequisite for clinical trust—models that collapse when moved between hospitals cannot be deployed safely.
@@ 3.5 Ensuring causal fairness in clinical algorithms
The RaVSNet framework demonstrates the clinical impact of causal fairness correction through explicit modelling of prescription bias [71]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how historical prescription patterns, physician preferences, and institutional protocols influence medication choices independently of therapeutic benefit. The system generates causal effect matrices, adjusting the probabilities of recommendations to reflect genuine therapeutic value rather than the tradition of prescribing. On MIMIC-III and MIMIC-IV datasets, RaVSNet achieved Jaccard scores of 0.573 and 0.497 respectively, representing 3–7% improvements over baseline methods, while the causal effect inference module specifically addresses confounding from historical prescribing patterns that may disadvantage underserved populations [71].
@@ 4.1.1 Mechanistic brain network analysis
Brain network analysis exemplifies the critical need for causal over correlational approaches. The human brain comprises approximately 86 billion neurons that form trillions of connections, creating one of the most complex networks in nature. Traditional graph-theoretical methods, while foundational, often fall short in capturing the high-dimensional and dynamic nature of brain connectivity [72]. The brain naturally forms a graph where regions of interest serve as nodes and functional connectivity defines edges, yet correlation-based analysis frequently identifies spurious connections driven by global brain states, arousal levels, or measurement artefacts rather than genuine neural communication pathways [53].
@@
The BPI-GNN framework advances psychiatric diagnosis through prototype learning that identifies biologically meaningful subtypes of psychiatric disorders [73]. Beyond discriminating patients from healthy controls, the model revealed subtypes with distinct clinical profiles. In major depressive disorder (REST-meta-MDD, n=1,604), identified subtypes exhibited distinct patterns on HAMD-17 clinical subscales including suicide risk, psychomotor retardation, and somatic symptoms, suggesting that network-based biomarkers could inform clinical characterisation. For schizophrenia (SRPBS, n=184), subtypes corresponded to differences in PANSS symptom profiles across positive, negative, and general psychopathology domains [73]. The model achieved superior performance against 11 popular brain network classification methods across three psychiatric datasets (REST-meta-MDD, ABIDE with n=1,064, and SRPBS), demonstrating both technical excellence and clinical relevance.
For epileptic seizure prediction, the Dynamic Instance-level Graph Learning Network (DIGLN) captures patient-specific causal relationships between intracranial electroencephalogram channels that evolve over time [65]. By learning both the organisation of epileptic networks and their temporal evolution, DIGLN demonstrated the ability to evaluate multiple overlapping patterns of brain functional change, particularly in connectivity between the visual and sensorimotor domains [74]. In 5-fold cross-validation on the Freiburg iEEG dataset (9 patients, 45 seizures), this patient-specific approach achieved 85.7% recall and 91.7% precision (F1: 87.0%), demonstrating the potential of dynamic graph learning for seizure prediction, though external validation on larger cohorts remains necessary.
@@ 4.1.2 Multi-omics cancer subtyping
Causal methodology transforms cancer classification by focusing on causally related genes rather than on all differentially expressed features. The CautionGCN framework employs a causal multi-head autoencoder for causal feature selection, isolating characteristic genes with genuine causal relationships to cancer subtypes [76]. This approach addresses a fundamental problem in cancer genomics, where thousands of genes show differential expression, but only a subset causally drives tumorigenesis. By constructing causal gene-gene graphs that capture driver relationships under assumptions of causal sufficiency and faithfulness (Table 2).
The gene-level resolution of MoCaGCN’s causal graph carries important validation implications where each identified causal driver can, in principle, be experimentally tested through CRISPR knockout or pharmacological inhibition of the specific gene product. This contrasts with pathway-level causal analyses where nodes aggregate multiple genes, creating ambiguity about which molecular targets should be perturbed to validate the discovered causal relationship.
The identified causal genes provide more than classification accuracy. The MOGONET framework takes advantage of both omics features and correlations among samples described by similarity networks, using the Correlation Discovery Network to explore cross-omics correlations in the label space [77]. This approach recognises that different omics layers may capture distinct aspects of cancer biology, with genomics revealing mutational drivers, transcriptomics showing pathway activation, and proteomics indicating functional consequences. By learning these multi-level representations jointly, MOGONET achieved superior performance in breast invasive carcinoma subtype classification while identifying layer-specific biomarkers.
@@ 4.2 Prognosis: predicting what interventions prevent
In a cohort of 593 patients with lung cancer surgery, KAMLN achieved micro-AUC of 0.664 for multi-label complication prediction by explicitly modelling causal dependencies between complications [78]. The knowledge-aware architecture enables earlier identification of causal chains, supporting prophylactic interventions before complications cascade. The interpretability of the model, showing not only risk, but also causal pathways, helped clinicians understand and trust the predictions, which proved crucial for clinical adoption. The explicit modelling of causal dependencies between complications has important equity implications beyond predictive performance. Patients in underserved communities, who often have more advanced disease states, benefit disproportionately from early identification of causal chains.